OpenStudy (lukecrayonz):

Find f(x) and g(x) so the function can be expressed as y = f(g(x)). y=(7/x^2)+10

OpenStudy (anonymous):

\[f(x)=x+10\] \[g(x)=7/x^2\]

OpenStudy (anonymous):

these can also be \[f(x)=7/x+10\] \[g(x)=x^2\]

OpenStudy (anonymous):

logic is g(x) is put inside f(x), break the equation to simpler variables e.g\[\sqrt{x^2+2x+2}\] \[f(x)=\sqrt{x}\] \[g(x)=x^2+2x+2\]